# Electroweak Monopole-Antimonopole Pair Production at LHC

### v2

@@ Abstract
One of the urgent issues in high energy physics is the experimental confirmation of the electroweak monopole predicted by the standard model, and currently MoEDAL at LHC is actively searching for the monopole. However, the present LHC cannot produce the monopole if the mass is bigger than 7 TeV, while the monopole mass is expected to be around $M_{W}/\alpha\simeq 11~{}\text{TeV}$. In this paper we discuss how LHC could circumbent this energy constraint and produce the monopole even when the mass becomes bigger than 7 TeV, based on the following ideas. First, in the topological production of the monopole the baby monopole mass at creation could be considerably smaller than the adolescent mass. Second, the binding energy of the monopole-antimonopole pair could effectively reduce the mass of the bound state. We discuss how these ideas can actually be realized at LHC to produce the monopole pairs. In particular, we argue that LHC could produce the baby electroweak monopoles whose mass could be around 5.3 TeV, smaller than the adolescent monopole mass around 11.0 TeV. Moreover, we show that LHC could produce the monopolium bound state with mass around 2.5 TeV, even when the total mass of the baby monopole-antimonopole pair becomes around 10.6 TeV. Our analysis could play an important role for MoEDAL experiment.
electroweak monopole, baby monopole mass, adolescent monopole mass, electroweak monopolium, Bohr radius of electroweak monopolium, mass of the electroweak monopolium, Rydburg energy of electroweak monopolium, monopolium production at LHC
@@ I Introduction
With the advent of the Dirac’s monopole the magnetic monopole has become an obsession in physics, experimentally as well as theoretically [1, 2, 3]. After the Dirac monopole we have had the Wu-Yang monopole, the ’t Hooft-Polyakov monopole, the grand unification monopole [4, 5, 6, 7, 8, 9, 10, 11]. But the electroweak monopole (also known as the Cho-Maison monopole), despite the initial skepticism on the existence of such monopole in the standard model [12, 13], stands out as the most realistic monopole that could exist in nature and could actually be detected [14, 15, 16].
Indeed the Dirac monopole in electrodynamics should transform to the electroweak monopole after the unification of the electromagnetic and weak interactions, and the Wu-Yang monopole in QCD is supposed to make the monopole condensation to confine the color. Moreover, the ’tHooft-Polyakov monopole exists only in an hypothetical theory, and the grand unification monopole which could have been amply produced at the grand unification scale in the early universe probably has become completely irrelevant at present universe after the inflation [17, 18].
@@
To detect the electroweak monopole at LHC, we need to recall basic facts about the monopole [14, 15, 23, 24, 25, 26, 27, 28]. First, this is the monopole which exists within (not beyond) the standard model as the electroweak generalization of the Dirac monopole, which can be viewed as a hybrid between Dirac and ’t Hooft-Polyakov monopoles. Second, the magnetic charge of the monopole is not $2\pi/e$ but $4\pi/e$, twice that of the Dirac monopole. This is because the period of the electromagnetic U(1) subgroup of the standard model is $4\pi$, since the electromagnetic U(1) (partly) comes from the U(1) subgroup of SU(2). Third, the mass of the monopole $M$ is of the order of several TeV, probably between 4 to 10 TeV. This is because the mass basically comes from the same Higgs mechanism which makes the W boson massive, except that here the coupling is magnetic (i.e., $4\pi/e$). This makes the monopole mass $1/\alpha$ times heavier than the W boson mass, around 11.0 TeV. Despite this, the size of the monopole is set by the weak boson masses. This is because the monopole solution has the Higgs and W boson dressing which has the exponential damping fixed by the weak boson masses.
@@
Because of these distinctive features, MoEDAL detector could identify the monopole without much difficulty, if LHC could produce it. However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV. This is problematic, because the monopole mass could turn out to be bigger than this, around 11.0 TeV. If this is the case, we may have little chance to detect the monopole at LHC. In this case we may have to try to detect the remnant monopoles at present universe produced in the early universe [27], or wait for the next upgrade of LHC.
@@
However, two observations could make the monopole signal at LHC non-zero even when the mass becomes over 7 TeV. First, the monopole has to be created in pairs at LHC. This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole, or the quantized magnetic flux of the monopole-antimonopole pair. In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass becomes bigger than 7 TeV.
@@
The second point is related to the electroweak monopole production mechanism at LHC. At present there are three contending monopole production mechanisms. Two popular ones in the high energy physics community are the Drell-Yan (and two photon fusion) process [19, 20] and the Schwinger mechanism [22, 29, 30]. The third one is the topological monopole production mechanism, in which the thermal fluctuation of the Higgs vacuum which takes place during the electroweak phase transition produces the monopole [27]. LHC could produce the monopole with this mechanism because the fireball of the p-p collision could re-create the electroweak phase transition in the early universe, and thus produces the monopoles topologically by the thermal fluctuation of the Higgs vacuum.
@@
A unique feature of this topological thermal fluctuation mechanism is that this could allow LHC to produce the electroweak monopole even when the mass becomes heavier than 7 TeV. To understand this point we first have to keep in mind that the monopole mass we have been talking about is the mass at the zero temperature. However, in this thermal production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [27]. This tells that we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass. So, from now on we will call the initial monopole mass (the monopole mass at birth) as the “baby” monopole mass and the zero temperature mass as the “adolescent” monopole mass.
[…]
As we will see, just after the phase transition the baby monopole mass could be as low as 1.0 TeV. This would have been impossible if LHC produces the monopole by Drell-Yan or Schwinger mechanism, where the concept of the temperature dependent monople mass does not exist. On the other hand, we emphasize that the monopole production in the early universe and the monopole production at LHC have important differences. First, the monopole production in the early universe stops around the Ginzburg temperature, while the monopole production at LHC stops at higher temperature. This is because at the present LHC the monopole mass can not exceed 7 TeV. So LHC has less monopole production time. Second, the size of the thermal radiation bath in the early universe is limited only by the horizon distance, but the size of the fireball at LHC in p-p collision is fixed by the proton size which is very small. Because of this, the monopole in the early universe has unlimited energy nearby but the monopole at LHC has very limited energy coming from the p-p fireball. These differences play important roles in the monopole production at LHC.
The existence of a monopole-antimonopole bound state, the metastable quantized magnetic flux made of Cho-Maison monopole-antimonopole pair, has recently been confirmed [31]. It has the expected pole separation length around $8.2/M_{W}$ and has the magnetic dipole moment roughly $8.6/M_{W}$. This extended size (length) of the magnetic flux suggests that the energy of the pair might be bigger than the sum of the monopole and antimonopole masses. This implies that LHC might need more energy to produce the magnetic flux made of the monopole-antimonopole pair than to produce the monopole and antimonopole separately.
This makes the monopolium bound state a more probable signal for the monopole production at LHC. The monopolium itself has been studied before for various reasons [32, 33, 34]. For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum to the order of hundred MeV, the mass scale of the low lying hadrons [34]. But the electroweak monopolium made of the Cho-Maison monopole pair has not been studied before.
Our study on the electroweak monopolium tells the followings. First, the naive atomic model of monopolium does not describe a realistic monopolium, because it has unrealistically big binding energy and too small size. This is because the magnetic Coulombic coupling of the monopole-antimonopole pair is too strong and the monopole mass is too big.
But we could avoid these difficulties and have a realistic electroweak monopolium modifying the magnetic potential at short distance. As we will see, in this modification we can show that the binding energy reduces the monopolium mass from 22.0 TeV to 5.3 TeV. This is precisely what we hoped for. With this we could conclude that LHC could produce the electroweak monopolium of the mass considerably lighter than 14 TeV without much difficulty. In this case we could try to detect the monopolium bound state experimentally detecting the decay modes.
@@
The paper is organized as follows. In Section II we discuss the electroweak monopole production mechanism at LHC. In Section III we review the monopole production in the electroweak phase transition in the early universe. In Section IV we discuss how LHC could produce the monopole in a similar manner, and argue that LHC could produce the baby electroweak monopole whose mass is smaller than the adolescent mass. In Section V we discuss the naïve Bohr model of the electroweak monopolium to see whether LHC could produce such electroweak monopolium. In Section VI we discuss a more realistic electroweak monopolium that could be produced at LHC. In Section VII we discuss the physical implications of our results.
II Electroweak Monopole Production Mechanisms at LHC
Before we discuss the electroweak monopole production at LHC, we review the standard model. Let us start from the (bosonic sector of the) Weinberg-Salam Lagrangian,
@@ II Electroweak Monopole Production Mechanisms at LHC
where $\phi$ is the Higgs doublet, $\vec{F}_{\mu\nu}$ and $G_{\mu\nu}$ are the gauge field strengths of $SU(2)$ and $U(1)_{Y}$ with the potentials ${\vec{A}}_{\mu}$ and $B_{\mu}$, and $g$ and $g^{\prime}$ are the corresponding coupling constants, and $\cal D_{\mu}$ is the covariant derivative of the $SU(2)$ and $U(1)_{Y}$ gauge groups. Expressing $\phi$ with the Higgs field $\rho$ and unit doublet $\xi$ by
@@
$F^{\prime}_{\mu\nu}=\partial_{\mu}A^{\prime}_{\nu}-\partial_{\nu}A^{\prime}_{\mu},$
$A^{\prime}_{\mu}=A_{\mu}+C_{\mu},~{}~{}~{}C_{\mu}=-\frac{2i}{g}\xi^{\dagger}\partial_{\mu}\xi$
@@
We emphasize that this is not the Weinberg-Salam Lagrangian in the unitary gauge. This is the gauge independent abelianization of the Weinberg-Salam Lagrangian. Notice that in this abelianization the Higgs doublet disappears completely, and the Higgs, W, and Z boson acquire the mass $M_{H}=\sqrt{\lambda}\rho_{0}$, $M_{W}=g\rho_{0}/2$, $M_{Z}=\sqrt{g^{2}+g^{\prime 2}}\rho_{0}/2$.
The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $\mathbb{C}P^{1}$ field. It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [14]. Although the Cho-Maison monopole has infinite energy classically, one could predict the mass to be around 4 to 11 TeV. Intuitively we could argue the mass to be $1/\alpha$ times heavier than the W boson mass, around 11 TeV. This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$). We could backup this arguement regularizing the Cho-Maison monopole, and estimate the mass to be around 4 to 10 TeV [23, 24, 25, 27, 28]. But in the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
To understand the electroweak monopole production at LHC we have to know the monopole production mechanism. In the Drell-Yan process the monopole-antimonopole pairs are produced by the electromagnetic interaction through the photon [19, 20]. But in the Schwinger mechanism a strong magnetic field creates the monopole-antimonopole pairs, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [22, 29, 30]. The justification for this comes from the assumption that a theory of monopole must have the dual symmetry, i.e., must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to $(g,-e)$. The Drell-Yan process is shown graphically in Fig. 1.
Notice, however, that something is strange with these pictures. For instance, in the Drell-Yan process the monopole pair is produced by the photon solely by the electromagnetic process. But the Weinberg-Salam Lagrangian has no interaction which can be described by the above Feynman diagram. Similarly, for the Schwinger mechanism the one-loop effective action of the Weinberg-Salam theory has no indication of the monopole pair production in strong magnetic background. Most importantly, the monopole is not an ordinary point particle but a topological particle, so that we need a change of topology to produce it. But it is not clear how the Drell-Yan process or the Schwinger mechanism could accommodate this topological nature of the monopole.
@@
The standard model provides another monopole production mechanism, the topological monopole production generated by the thermal fluctuation of the Higgs vacuum which induces the change of topology, which is the electroweak monopole production mechanism in the early universe [27]. In this picture the monopoles are produced in the course of the electroweak phase transition by thermal fluctuations of the Higgs vacuum, which takes place after the phase transition and stops at the Ginzburg temperature [27]. This is shown in Fig. 2, which should be contrasted with Fig. 1.
A remarkable feature of this mechanism is that the monopole production is not controlled by any fundamental interaction, but by thermal fluctuation which does not involve any fundamental constant. Perhaps more importantly, in this thermal fluctuation the monopole mass at creation (the baby monopole mass) could be considerably smaller than the adolescent monopole mass. This is because in this thermal production, the monopole mass depends on the temperature at creation. As we will see in the following this could have a deep impact on the monopole production at LHC.
If so, what is the the monopole production mechanism at LHC? We do not know yet. On the other hand, it has often been claimed that LHC should reproduce the early universe. If this is true, the monopole production mechanism at LHC could be topological, just like the electroweak monopole production the early universe. In this paper we will adopt this view and discuss how LHC could produce the electroweak monopole. For this we have to understand the electroweak monopole production in the early universe.

### v4

@@ Abstract
The monopole production at LHC crucially depends on the monopole production mechanism. We show that, if the monopole production mechanism at LHC is the thermal fluctuation of the Higgs vacuum as the early universe did, it is practically impossible for LHC to produce the monopole, even with the future FCC energy upgrade. This is because the temperature of the p-p fireball is simply too low to produce the monopole. But if the monopole production mechanism is the Drell-Yan and/or Schwinger mechanism, the 14 TeV LHC could produce the monopole in the form of the monopolium of mass around 5,7 TeV even when the monopole mass becomes 11 TeV. Or, it could produce the monopole when the mass is less than 7 TeV. Our result tells that, if the monopole production mechanism at LHC becomes the thermal fluctuation, searching for the remnant monopoles produced in the early universe could be the only way to detect the electroweak monopole. We discuss the physical implications of our result.
electroweak monopole, monopole production mechanism at LHC, Drell-Yan process, Schwiner mechanism, thermal fluctuation of Higgs vacuum, Ginzburg temperature, baby monopole mass, adolescent monopole mass, electroweak monopolium, Bohr radius of electroweak monopolium, mass of the electroweak monopolium, Rydburg energy of electroweak monopolium, monopolium production at LHC
@@ I Introduction
With the advent of the Dirac’s monopole the magnetic monopole has become an obsession in physics, experimentally as well as theoretically [1, 2]. After the Dirac monopole we have had the Wu-Yang monopole, the ’t Hooft-Polyakov monopole, and the grand unification monopole [3, 4, 5]. But the electroweak (Cho-Maison) monopole, stands out as the most realistic monopole that could exist in nature and could actually be detected [6, 7].
Indeed the Dirac monopole in electrodynamics should transform to the electroweak monopole after the unification of the electromagnetic and weak interactions, and the Wu-Yang monopole in QCD is supposed to make the monopole condensation to confine the color. Moreover, the ’tHooft-Polyakov monopole exists only in an hypothetical theory, and the grand unification monopole which could have been amply produced at the grand unification scale in the early universe probably has become completely irrelevant at present universe after the inflation.
@@
To detect the electroweak monopole at LHC, we need to recall basic facts about the monopole [6, 7, 12, 13, 14, 15, 16]. First, this is the monopole which exists within (not beyond) the standard model as the electroweak generalization of the Dirac monopole, which can be viewed as a hybrid between Dirac and ’t Hooft-Polyakov monopoles. Second, the magnetic charge of the monopole is not $2\pi/e$ but $4\pi/e$, twice that of the Dirac monopole. This is because the period of the electromagnetic U(1) subgroup of the standard model becomes $4\pi$. Third, the mass of the monopole $M$ is of the order of 10 TeV. This is because the mass basically comes from the same Higgs mechanism which makes the W boson massive, except that here the magnetic coupling makes the monopole mass $1/\alpha$ times heavier than the W boson mass, around 11.0 TeV [17]. Despite this, the size of the monopole is set by the W boson mass, because the monopole solution has the Higgs and W boson dressing which fixes the size by the W boson mass.
@@
Because of these distinctive features, MoEDAL detector could identify the monopole without much difficulty, if LHC could produce it. However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV. This is problematic, because the monopole mass could turn out to be bigger than this. If this is so, we may have little chance to detect the monopole at LHC, and may have to try to detect the remnant monopoles at present universe produced in the early universe [15, 18].
@@
However, two observations could make the monopole production at LHC possible even when the mass becomes over 7 TeV. First, the monopole has to be created in pairs at LHC. This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole. In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass becomes bigger than 7 TeV.
@@
The second point is related to the electroweak monopole production mechanism at LHC. At present there are three contending monopole production mechanisms. Two popular ones in the high energy physics community are the Drell-Yan (and two photon fusion) process [8, 9] and the Schwinger mechanism [10, 19, 20]. The third one is the topological monopole production mechanism, in which the thermal fluctuation of the Higgs vacuum produces the monopole as the early universe did [15, 18]. LHC could produce the monopole with this mechanism because the fireball of the p-p collision could reproduce the early universe.
@@
A unique feature of this topological thermal fluctuation mechanism is that in this thermal production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [15, 18]. In this case there is a possibility that LHC could produce the electroweak monopole even when the zero temperature mass becomes heavier than 7 TeV.
The purpose of this paper is to discuss how LHC can accommodate the above ideas and produce the electroweak monopoles even when the monopole mass exceeds 7 TeV. We show that, if the monopole production mechanism at LHC is the Drell-Yan process (and/or Schwinger mechanism), the present 14 TeV LHC could produce the monopole if the mass is less than 7 TeV. Or, it could produce the monopole in the form of the monopolium bound states of mass around 5.7 TeV even when the monopole mass becomes 11 TeV. This is because the binding energy of the monopoliun could reduce the monopolium mass by 16.3 TeV. Indeed we argue that the LHC can produce the monopolium for any reasonable monopole mass, as far as the mass does not exceed 29.8 TeV. However, if the monopole production mechanism at LHC becomes the thermal fluctuation of the Higgs vacuum, it has practically no chance to produce the monopole even with the future FCC energy upgrade. This is because the temperature of the p-p fireball at LHC simply becomes too low to produce the monopole thermally. This makes the monopolium bound state a most probable signal for the monopole production at LHC.
@@
The paper is organized as follows. In Section II we discussthe electroweak monopole production mechanism at LHC. In Section III we review the monopole production in the electroweak phase transition in the early universe. In Section IV we discuss if LHC could produce the monopole in a similar manner, and argue that this is practically impossible because the temperature of the p-p fireball is too low. In Section V we discuss the naive Bohr model of the electroweak monopolium to see whether LHC could produce such electroweak monopolium. In Section VI we discuss a more realistic electroweak monopolium that could be produced at LHC. In Section VII we discuss the physical impications of our results.
II Electroweak Monopole Production Mechanism at LHC
Consider the (bosonic sector of the) Weinberg-Salam Lagrangian,
@@ II Electroweak Monopole Production Mechanism at LHC
where $\phi$ is the Higgs doublet, $\vec{F}_{\mu\nu}$ and $G_{\mu\nu}$ with potentials ${\vec{A}}_{\mu}$ and $B_{\mu}$ are the gauge fields of $SU(2)$ and $U(1)_{Y}$, $\cal D_{\mu}$ is the covariant derivative, and $g$ and $g^{\prime}$ are the corresponding coupling constants. Expressing $\phi$ with the Higgs field $\rho$ and unit doublet $\xi$ by
@@
$F^{\prime}_{\mu\nu}=\partial_{\mu}A^{\prime}_{\nu}-\partial_{\nu}A^{\prime}_{\mu},~{}~{}A^{\prime}_{\mu}=A_{\mu}+C_{\mu},~{}~{}C_{\mu}=-\frac{2i}{g}\xi^{\dagger}\partial_{\mu}\xi$
@@
Notice that here we have defined $A_{\mu}^{(em)}$ and $Z_{\mu}$ gauge independently, without choosing the unitary gauge.
@@
We emphasize that this is not the Weinberg-Salam Lagrangian in the unitary gauge. This is the gauge independent abelianization of the Weinberg-Salam Lagrangian. Notice that here the Higgs doublet disappears completely, and the Higgs, W, and Z bosons acquire the mass $M_{H}=\sqrt{\lambda}\rho_{0}$, $M_{W}=g\rho_{0}/2$, $M_{Z}=\sqrt{g^{2}+g^{\prime 2}}\rho_{0}/2$, without any spontaneous symmetry breaking.
@@
The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $\mathbb{C}P^{1}$ field. It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [6, 7]. Although the Cho-Maison monopole has infinite energy classically, one could predict the mass to be around 4 to 11 TeV. Intuitively we could argue the mass to be $1/\alpha$ times bigger than the W boson mass, around 11 TeV [17]. This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$). We could backup this arguement regularizing the Cho-Maison monopole, and estimate the mass to be around 4 to 10 TeV [12, 13, 14, 15, 16]. But in the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
To understand the electroweak monopole production at LHC we have to know the monopole production mechanism at LHC. In the Drell-Yan process the monopole-antimonopole are produced in pairs by the electromagnetic interaction through the photon [8, 9]. But in the Schwinger mechanism a strong magnetic field is supposed to create the monopole-antimonopole pair, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [10, 19]. The justification for this comes from the asumption that a theory of monopole must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to $(g,-e)$. The Drell-Yan process is shown graphically in Fig. 1.
But in the topological monopole production, the thermal fluctuation of the Higgs vacuum induces the change of topology and produce the monopole, which is exactly the electroweak monopole production mechanism in the early universe [15]. In this picture the monopoles are produced after the phase transition and stops at the Ginzburg temperature [15]. The topological monopole production mecanism is shown graphically in Fig. 2 for comparison.
Notice that, in the Drell-Yan process the monopole pair is produced by the photon solely by the electromagnetic process. But the Weinberg-Salam Lagrangian perturbatively has no interaction which can be described by the above Feynman diagram. Similarly, for the Schwinger mechanism the one-loop effective action of the Weinberg-Salam theory has no indication of the monopole pair production in strong magnetic background. Most importantly, we need the change of topology to produce the monopole, but this topological nature of the monopole has no role at all in these mechanisms.
In comparison the topology plays an essential role in the thermal production of the monopole, because here the thermal fluctuation of the Higgs vacuum is precisely what we need to induce the change of topology. As a result the monopole production in this mechanism is not controlled by any fundamental interaction or fundamental constant. Perhaps more importantly, in this thermal fluctuation the monopole mass depends on the temperature at creation, the monopole mass at creation could be considerably smaller than the monopole mass at zero temperature. In this case we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass. So, from now on we call the initial monopole mass (the monopole mass at birth) as the “baby” (or “infant”) monopole mass and the zero temperature mass as the “adolescent” monopole mass.
It should be emphasized that the above monopole production mechanisms at LHC are the logical possibilities, not the realistic possibilities. For example, the Schwinger mechanism is certainly a logical possibility. However, it is not clear at all if (and how) the p-p fireball at LHC could create a strong magnetic background which can actually produce the monopole-antimonopole pair.
If so, what is the the monopole production mechanism at LHC? We do not know yet. On the other hand, it has often been claimed that LHC could reproduce the early universe. If this is true, the monopole production mechanism at LHC could be topological, just like the electroweak monopole production in the early universe. In the following we discuss if LHC could produce the electroweak monopole thermally.